# Extreme Compression of Large Language Models via Additive Quantization

### 11 Jan 2024 v1

@@ 3.4 Fine-tuning for intra-layer cohesion
7B 16.00 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 4.044 5.21 6.74 67.32 77.69 56.10 68.90 40.27 62.06
GPTQ 4.000 5.49 7.20 67.17 76.82 55.17 67.42 40.02 61.32
SpQR 3.980 5.28 6.87 66.54 77.42 55.97 68.14 39.85 61.58
QuIP 4.000 - - - - - - - -
QuIP# 4.020 5.29 6.86 66.85 77.91 55.78 68.06 39.68 61.66
13B 16.00 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 3.938 4.64 6.14 70.40 78.56 59.23 72.22 43.94 64.87
GPTQ 4 4.78 6.34 68.51 77.48 58.57 69.82 42.58 63.39
SpQR 3.980 4.69 6.20 68.82 78.13 59.11 72.90 44.62 64.72
QuIP 4.00 4.76 6.29 69.69 79.00 58.91 73.27 44.88 65.15
QuIP# 4.011 4.68 6.20 69.38 77.91 58.86 73.74 44.63 64.90
70B 16.00 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 4.07 3.17 5.01 76.90 81.41 63.64 78.09 51.00 70.21
GPTQ 4.00 3.34 5.15 75.45 81.56 63.24 76.56 49.57 69.28
SpQR 3.97 3.24 5.07 76.48 81.12 63.70 76.05 50.34 69.54
QuIP 4.00 3.58 5.38 76.01 80.25 61.97 74.28 47.01 67.90
QuIP# 4.01 3.22 5.05 76.80 81.45 63.51 78.37 50.85 70.20
4 Experiments
We evaluate the AQLM algorithm in typical scenarios for post-training quantization of modern LLMs. Our evaluation is focused on the Llama 2 model family since it is a popular backbone for fine-tuned models or general LLM applications, e.g. (Dettmers et al., 2023a). We organize the experiments as follows: in Section 4.1 we evaluate the full AQ procedure for various Llama 2 models and quantization bit widths, and Section 4.2 presents an ablation analysis for individual AQ components and implementation details.
4.1 Llama 2 compression
To evaluate the quality of AQ-quantized models, we report perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) validation sets. We also measure zero-shot accuracy on five tasks: WinoGrande (Sakaguchi et al., 2021), PiQA (Tata & Patel, 2003), HellaSwag (Zellers et al., 2019), ARC-easy and ARC-challenge (Clark et al., 2018) via the LM Evaluation Harness (Gao et al., 2021). We run all evaluations with the same hyperparameters as in GPTQ (Frantar et al., 2022a), except for the increased number of training calibration sequences.
We consider three main targets in terms of compression range: 2-2.1 bits, 3-3.1 bits, and 4-4.1 bits per model parameter. The full number of bits for each algorithm includes scales, codebooks, zero points, but not the (non-quantized) word embeddings and logits. We compare AQ against GPTQ for 3&4 bits (Frantar et al., 2022a), SpQR for 3&4 bits (Dettmers et al., 2023b), QuIP in 2,3 & 4 bits (Chee et al., 2023) and QUIP# for 2&4 bits (Tseng et al., ). While GPTQ and SpQR technically support 2-bit quantization, these algorithms were designed for 3+ bits per parameter and perform poorly in terms of accuracy in the 2-2.1 bit range. For QuIP, we omit results for the 7B model, as we could not achieve competitive performance in this one scenario using the available [^4] Unfortunately, there is no official implementation of the original QuIP (non-#) for the Llama 2 model, see https://github.com/Cornell-RelaxML/QuIP/issues/2. implementations. For QuIP#, we focus on 2 and 4 bit because the available implementation does not yet support 3-bit compression. While there are several ways to modify QuIP to support 3 bits, it is unclear which way is more natural. We calibrate each quantization algorithm using the same portion of RedPajama calibration dataset (Computer, 2023), with a sequence length of 4096.
We report the results in Tables 1 2 and 3 (for 2, 3 and 4 bits respectively), both in terms of perplexity and zero-shot accuracy. To summarize, AQLM outperforms the previous best PTQ algorithms across all settings, but the margin of improvement varies significantly. For example, we observe the highest gains in perplexity in the “extreme” 2-2.1 bits per parameter range. One possible explanation for this is that layerwise fine-tuning (defined in Section 3.4) gives the most benefit in regions with lower compression bit width because these regions have higher error in the first place. Conversely, we observe that 3 and 4-bit quantizers are already fairly close to unquantized (FP16) model perplexity.
@@ 4.2 Ablation analysis
The AQLM algorithm makes several design choices that need to be validated separately: the initialization, the alternating optimization, the fine-tuning protocol, and the choice of hyperparameters. Here, we study how each of these components affect the overall algorithm.
@@
Initialization. As we discuss in Section 3, we initialize AQLM with residual K-means to obtain a good initial guess for both codes and codebooks. That is, we run K-means for the weight matrix, then subtract the nearest cluster from each weight, and run K-means again M times. A simple baseline would be to initialize all codes uniformly at random. We compare the two initialization strategies for the problem of quantizing a single linear layer within Llama 2 70B model to 3 bits per parameter. We quantize groups of 8 consecutive weights using 2 codebooks, 12 bit each. Each codebook contains ${2^{12}}$ learnable values. As we can see in Figure 4, AQLM with K-means initialization needs significantly fewer training iterations to achieve the desired loss. The difference is so drastic that we expect that running AQLM with a random initialization would require extremely high runtimes to accurately quantize the largest models.
@@
Number of samples. Finally, we verify our choice of calibration hyperparameters. Traditionally, most PTQ algorithms use several hundred calibration sequences (e.g. Frantar et al. (2022a) has 128). In our experiments, we evaluate both AQLM and baselines with additional calibration data. Our original motivation for that was to avoid potential overfitting when fine-tuning entire transformer blocks. To test this assumption, we run our algorithm with different calibration set sizes, varying from 128 to 4096 sequences. For each size, we report the average perplexity on WikiText2 over 3 runs, along with standard deviations. The results in Table 5 demonstrate that increasing the number of samples leads to gradual reduction in perplexity with seemingly diminishing returns. Since the perplexity is still monotonically improving from 128 to 4096 samples, it is possible that larger sample sizes would yield further improvements.
@@
5 Conclusion and Future Work
We presented AQLM, a variant of additive quantization (AQ) targeted to accurate LLM compression. Specifically, AQLM adapts AQ to the layer-wise quantization problem by making it instance-aware, by taking the layer input distributions into account into the codebook optimization. Moreover, we combine this approach with a block fine-tuning approach, which allows us to further reduce quantization error across layers. To our knowledge, AQLM improves the current state-of-the-art results for LLM quantization in the regime of 2 and 3 bits per weight.
In terms of limitations, AQLM is more computationally-expensive relative to existing direct post-training quantization methods, such as RTN or GPTQ. Moreover, as detailed in the experiments, we have found it to be sensitive to the choice of parameters, such as the number of codebooks, or the initialization of centroids. Further, since it uses a more sophisticated encoding, its implementation will be more complex than direct quantization methods.
In future work, we plan to provide an efficient GPU implementation of AQLM. Further, we intend to conduct a more detailed analysis of the importance of specific parameters on the quality of compression, as well as end-to-end joint fine-tuning of codebook parameters, across the entire model.
6 Acknowledgements
Authors would like to thank Ruslan Svirschevski for his help in solving technical issues with AQLM and baselines. We also thank Tim Dettmers for helpful discussions on the structure of weights in modern LLMs and size-accuracy trade-offs. Finally, authors would like to thank the communities of ML enthusiasts known as LocalLLaMA [^5] https://www.reddit.com/r/LocalLLaMA/ and Petals community on discord [^6] https://github.com/bigscience-workshop/petals/ for the crowd wisdom about running LLMs on consumer devices.

### 06 Feb 2024 v2

@@ 4.1 Compression quality for modern LLMs
8x7B 16 3.46 5.02 75.45 82.37 64.65 83.38 55.80 72.33
AQLM 1.98 4.61 5.75 73.64 79.27 57.91 78.96 48.63 67.68
QuIP# 2.01 4.75 5.89 71.11 79.05 58.23 77.57 45.73 66.34
Mixtral quantization.
Table 3 presents results on the Mixtral MoE-type model, comparing against QuIP# at 2-bits. (See Appendix E.1 for full results.) AQLM outperforms QuIP# in this case as well. Although the margins are lower compared to Llama 2 models, they are still significant for “harder” tasks, such as Arc Challenge (+3 points).
Pareto optimality of AQLM. The significant error improvements raise the question of choosing the “optimal” model variant to maximize accuracy within a certain memory budget. For this, we follow Dettmers & Zettlemoyer (2022): a quantized model is said to be Pareto-optimal if it maximizes accuracy at the same or lower total size (bytes). Despite rapid progress, prior art methods are not Pareto-optimal at 2-bits: for instance, the previous best 2-bit Llama 2 13B (QuIP#, Table 1) achieves Wiki2 PPL of 6.06, but one can get much lower 5.21 PPL by using a 7B model with 4-bit quantization, which is smaller (see Appendix Table 10).
AQLM compression to strictly 2 bits for the same model is also below Pareto-optimality, as it is outperformed by 4-bit AQLM compression for Llama 2 7B (5.21 vs 5.65). To find the Pareto-optimal quantization bitwidth, we run experiments between 2-3 bits per parameter and report them in Table 1, below horizontal bars. Thus, the Pareto-optimal bitwidth for AQLM appears to be around 2.5 bits per parameter (Table 1), at which point we are comparable to 5-bit AQLM for Llama 2 7B (Appendix Table 10). In turn, the 2.76-bit AQLM on 13B outperforms the uncompressed 7B model. As such, AQLM is the first algorithm to achieve Pareto-optimality at less than 3 bits per parameter.
@@ 4.2 Ablation analysis
In Appendix D, we examine key design choices regarding initialization, alternating optimization, the impact of the fine-tuning protocol, distribution of codebooks vs groups, as well as other hyper-parameters. In brief, we first find that the residual K-means initialization is critical for fast algorithm convergence: when compared with random initialization, it needs significantly fewer training iterations. Second, to validate our calibration fine-tuning procedure, we compare it against 1) no fine-tuning, 2) fine-tuning only of non-linear layers (e.g. RMSNorm) but not of codebook parameters, and 3) fine-tuning only the codebooks (but not other layers). The results, presented in full in Appendix D, show that fine-tuning the codebook parameters has the highest impact on accuracy, by far, while fine-tuning the RMSNorm only has minor impact. This validates our choice of leveraging the calibration set for learned codebooks.
Further, we observe that increasing the number of sample sequences in the range 128 to 4096 leads to a gradual PPL improvement, but with diminishing returns. In this respect, AQLM benefits more from larger calibration sets (similarly to QuIP#), as opposed to direct methods like GPTQ which saturate accuracy at around 256 input sequences. Finally, we investigate various options for investing a given bit budget, comparing e.g. longer codes (e.g. 1x15) vs multiple codebooks with shorter codes (e.g. 2x8).
4.3 Inference Speed
Although our primary objective is to maximize accuracy for a given model size, AQLM can also be practical in terms of inference latency. To demonstrate this, we implemented efficient GPU and CPU kernels for a few hardware-friendly configurations of AQLM. The results can be found in Table 4. For GPU inference, we targeted quantized Llama 2 models with 16-bit codebooks, corresponding to 2.07 bits for Llama 2 70B, 2.18 bits for 13B, and 2.29 bits for 7B models (see Table 1), as well as a 2x8-bit codebook model with perplexity 7.98 on Wiki2. For each model we benchmark the matrix-vector multiplication subroutine performance on a standard layer. The results show that AQLM can execute at speeds comparable to or better than FP16. End-to-end generative numbers on a preliminary HuggingFace integration can be found in Appendix H: for instance, we can achieve $\sim$ 6 tokens/s on Llama 2 70B in this setting. We observe that multiple smaller codebooks allow efficient GPU cache utilization, leading to greater speedup, at the price of slightly lower accuracy.
Table 4: Speed of the FP16 gate_proj layer matrix-vector multiplication in PyTorch, and relative AQLM speedups.
Llama 2 7B 13B 70B
2 bit speedup over FP16 on Nvidia RTX 3090 GPU
Original (float16) 129 $\mu$s 190 $\mu$s 578 $\mu$s
AQLM (Table 1) x1.31 x1.20 x1.20
AQLM (2×2$\times$2 ×8-bit) x1.57 x1.82 x3.05
2 bit speedup over FP32 on Intel i9 CPU, 8 cores
Original (float32) 1.83 ms 3.12 ms 11.31 ms
AQLM (2×2$\times$2 ×8-bit) x2.75 x3.54 x3.69
AQLM (4×4$\times$4 ×8-bit) x2.55 x3.02 x4.07
AQLM (8×8$\times$8 ×8-bit) x2.29 x2.68 x4.03
Next, we explore how to leverage AQLM to accelerate CPU inference. As discussed in Section 2.2, additive quantization can compute dot products efficiently if the codebook size is small. One way to achieve it for AQLM is to replace each 16-bit codebook with a number of smaller 8-bit ones. This leads to higher quantization error, but still outperforms the baselines in terms of accuracy (see Appendix Table 9). The results in Table 4 show that this also allows for up to 4x faster inference relative to FP32 on CPU.
5 Conclusion and Future Work
We presented AQLM, a new form of additive quantization (AQ) targeted to LLM compression, which significantly improved the state-of-the-art results for LLM quantization in the regime of 2 and 3 bits per weight. In terms of limitations, AQLM is more computationally-expensive than direct post-training quantization methods, such as RTN or GPTQ, specifically because of the use of a more complex coding representation. Yet, despite the more sophisticated encoding and decoding, we have shown AQLM lends itself to efficient implementation on both CPU and GPU. Overall, we find it remarkable that, using AQLM, massive LLMs can be executed accurately and efficiently using little memory.
6 Acknowledgements
Authors would like to thank Ruslan Svirschevski for his help in solving technical issues with AQLM and baselines. We also thank Tim Dettmers for helpful discussions on the structure of weights in modern LLMs and size-accuracy trade-offs.The authors would also like to thank Daniil Pavlov for his assistance with CPU benchmarking. Finally, authors would like to thank the communities of ML enthusiasts known as LocalLLaMA [^3] https://www.reddit.com/r/LocalLLaMA/ and Petals community on discord [^4] https://github.com/bigscience-workshop/petals/ for the crowd wisdom about running LLMs on consumer devices.
Appendix A Full model fine-tuning
The block-wise finetuning procedure, introduced in 3.4, considerably improves performance of compressed models. However, block-wise finetuning optimizes the loss only at the level of a current transformer block and is agnostic of the actual task of interest. To minimize the target loss, one can run backpropagation through the whole model and optimize all trainable parameters to directly optimize the end objective.
This allows to search for globally optimal parameters, as opposed to sequentially selected ones, during block-wise finetuning.
One can minimize the error between the quantized model and the floating-point model on some calibration set. The parameters being optimized (namely the codebooks, scales and the non-quantized parameters) typically constitute a small fraction of the total number of parameters in the original model. Therefore, the proposed distillation method is an instance of parameter-efficient finetuning.
Standard Knowledge distillation (Hinton et al., 2015) optimizes the KL-divergence between the outputs of teacher and student models. However, according to the literature, (Sun et al., 2019; Kurtic et al., 2023) transfer of intermediate representation facilitates faster convergence and enhances performance. Hence, we decided to perform knowledge transfer between some intermediate activations in addition to the model output. Specifically, we take a subset of transformer blocks $\{l_{0},l_{1}\ldots,l_{K-1}\},l_{i}\in[0,L-1]$, where $L$ is the number of transformers block in the model, and optimize the following objective:
$\sum_{i=0}^{K-1}\mathcal{L}(x_{l_{i}}^{(s)},x_{l_{i}}^{(t)})$ (9)
Where $x_{l_{i}}^{(s)}$ and $x_{l_{i}}^{(t)}$ are the student and teacher output activations respectively at block $l_{i}$ and $\mathcal{L}(*,*)$ is some loss function. We considered the following options:
  • $L_{2}$ loss: $\|x-y\|_{2}^{2}$
  • Normalized $L_{2}$ loss: $\|x-y\|_{2}^{2}/\|y\|_{2}^{2}$
  • Cosine distance: $\frac{1}{2}\left(1-\frac{(x,y)}{\|x\|\|y\|}\right)$
We tune all models for a single epoch on the same calibration set used for additive quantization and block finetuning. We observe that longer training leads to overfitting. Learning rate of $\eta=1e-5$ and batch size of $32$ with Adam optimizer are used in full model finetuning experiments. We find that simple $L_{2}$ loss outperforms the alternatives (Table 5), and that the number of intermediate activations used in the loss function barely affects the final performance (Table 6).
Table 5: Ablation on the choice of loss for Knowledge distillation.
Loss Wiki2normal-$\downarrow$ C4normal-$\downarrow$
$L_{2}$ $\mathbf{6.57}$ $\mathbf{8.48}$
Normalized $L_{2}$ 6.63 8.57
Cosine distance 6.65 8.59
Table 6: Ablation on the choice of blocks subset for Knowledge distillation. Empty set means no knowledge distillation applied.
Blocks Wiki2normal-$\downarrow$ C4normal-$\downarrow$
[…]
$\{31\}$ $\mathbf{6.52}$ 8.51
$\{15,31\}$ 6.57 $\mathbf{8.48}$
$\{7,15,23,31\}$ 6.55 8.49
$\{3,7,11,15,19,23,27,31\}$ 6.57 8.54
Overall, the improvement due to knowledge distillation is not very significant while increasing the complexity of the quantization pipeline, therefore, we decided to not include KD in our main setup.
Appendix B Code reproducibility
We share the code for our method in the GitHub repository https://github.com/vahe1994/AQLM. The hyperparameters for our experimental setup are discussed in Appendix C.
Appendix C Experimental Configurations
Hardware. In all of our experiments, we used either Nvidia A100 or H100. The number of GPUs varied from 1 to 8. We used activation offloading to lower pick memory usage. To evaluate inference speed on GPU we used consumer-grade GPU Nvidia 3090 and for CPU setup we used Intel core i9 13900k.
Calibration set. All methods were calibrated on a slice of RedPajama-v1 dataset (Computer, 2023) for both Llama and Mistral/Mixtral family models. We used the same context length as models were trained on, for Llama-2 4096 and for Mistral/Mixtral 8192.
For Llama-2 experiments, we used 8M tokens as a calibration set for SpQR, GPTQ, and AQLM. Quip, however, was calibrated on 4M tokens due to OOM errors when trying to use more samples. Taking into account the fact that after 2M tokens improvement of methods results is fairly small we chose to report these numbers as is. For Quip#, we used Llama-2 and Mistral’s quantized models provided by authors in their GitHub repository. To the best of our knowledge, they used 6k samples for calibration with a context length of 4096/8192.
For Mixtral we calibrated both our method and QUIP# on 8M tokens with context length 8192.
Hyperparameters.
For GPTQ for both 3 and 4 bits we used a standard set of parameters without grouping and with permutation order act_order.
SpQR method was evaluated with base 2 and 3 bit-width with group size of 16 and 3 bits for zeros and scales. Outliers rate was chosen such that average bit will be close to 3 and 4 bits respectively.
Quip was adapted to work on the Llama family and was calibrated with 1024 samples and 4096 context length.
Quip# For Llama-2 and Mistral models we used the officially published quantized models. For Mixtral we adapted the code to work with the model’s architecture and quantized it with the recommended set of parameters. For both AQLM and QuIP#, we don’t quantize gate linear layer in Mixtral, because it contains relatively small amount of paramters and have severe impact on performance.
AQLM For to get 2, 3, 4 bits: we used 1 codebook size of $2^{15}$ or $2^{16}$, with groups of 8 for 2 bits. For 3 bits we used 2 codebooks size of $2^{12}$ with groups of 8. Finally for 4 bits we used 2 codebooks size of $2^{15}$ or $2^{16}$ with groups of 8.
Both for finetuning 3.4 and codebooks update 3.3 we used Adam optimizer (Kingma & Ba, 2015) with learning rate of $10^{-4}$, $\beta_{1}=0.90$ and $\beta_{2}=0.95$. We used early stopping both for the finetuning phase and for the codebook optimization phase, by stopping when the least square error not decreasing more than some threshold. In our experiments the threshold varies between $10^{-2}$ and $10^{-3}$.
Appendix D Ablation analysis
@@
The AQLM algorithm makes several design choices that need to be validated separately: initialization, alternating optimization, the fine-tuning protocol, and the choice of hyperparameters. Here, we study how each of these components affect results.
@@
Initialization. As discussed in Section 3, we initialize AQLM with residual K-means to obtain a good initial guess for both codes and codebooks. That is, we run K-means for the weight matrix, then subtract the nearest cluster from each weight, and run K-means again M times. A simple baseline would be to initialize all codes uniformly at random. We compare the two initialization strategies for the problem of quantizing a single linear layer within Llama 2 70B model to 3 bits per parameter. We quantize groups of 8 consecutive weights using 2 codebooks, 12 bit each. Each codebook contains ${2^{12}}$ learnable values. As we can see in Figure 4, AQLM with K-means initialization needs significantly fewer training iterations to achieve the desired loss. The difference is so drastic that we expect that running AQLM with a random initialization would require extremely high runtimes to accurately quantize the largest models.
@@
Number of samples. We verify our choice of calibration hyperparameters. Traditionally, most PTQ algorithms use several hundred calibration sequences (e.g. Frantar et al. (2022a) has 128). In our experiments, we evaluate both AQLM and baselines with additional calibration data. Our original motivation for that was to avoid potential overfitting when fine-tuning entire transformer blocks. To test this assumption, we run our algorithm with different calibration set sizes, varying from 128 to 4096 sequences. For each size, we report the average perplexity on WikiText2 over 3 runs, along with standard deviations. The results in Table 8 demonstrate that increasing the number of samples leads to gradual reduction in perplexity with seemingly diminishing returns. Since the perplexity is still monotonically improving from 128 to 4096 samples, it is possible that larger sample sizes would yield further improvements.
Number of codebooks vs groups. Finally, we conducted an additional set of experiments on LLama-2 7B models to see perplexity dependence on simultaneous change of both codebooks and groups keeping compression rate fixed to 2bits. 9
We can see that the simultaneous increase in both the number of codebooks and groups led to a decrease in average perplexity and stopped when the number of groups reached 64.
@@
Table 9: Wikitext2 PPL as a function of from groups and number of codebook for Llama 2 (7B) quantized with approximately 2 bits quantization.
# of codebooks Groups size Average PPL
1 4 41.79
2 8 11.96
4 16 8.53
8 32 7.83
15 64 7.86
Appendix E Additional experiments
In this section we report additional experimental results for Mixtral(Jiang et al., 2024) , Mistral7B(Jiang et al., 2023) and LLama-2 model.
Table 10: Evaluation of quantized Llama 2 models for 4+ bits per parameter. The table reports perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
7B 16 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 4.04 5.21 6.75 67.32 78.24 55.99 70.16 41.04 62.55
GPTQ 4.00 5.49 7.20 68.19 76.61 55.44 66.20 36.77 60.64
SpQR 3.98 5.28 6.87 66.93 78.35 56.10 69.11 39.68 62.17
QuIP# 4.02 5.29 6.86 66.85 77.91 55.78 68.06 39.68 61.66
AQLM 5.02 5.16 6.68 67.40 78.29 56.53 68.94 39.93 62.22
13B 16 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 3.94 4.65 6.14 69.85 78.35 59.27 73.32 44.80 65.12
GPTQ 4 4.78 6.34 70.01 77.75 58.67 70.45 42.49 63.87
SpQR 3.98 4.69 6.20 69.69 78.45 59.25 71.21 44.52 64.42
QuIP 4.00 4.76 6.29 69.69 79.00 58.91 73.27 44.88 65.15
QuIP# 4.01 4.68 6.20 69.38 77.91 58.86 73.74 44.63 64.90
70B 16 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 4.14 3.19 5.03 76.48 81.50 63.69 77.31 50.68 69.93
GPTQ 4.00 3.35 5.15 75.61 81.23 63.47 76.81 49.15 69.25
SpQR 3.97 3.25 5.07 76.01 81.28 63.71 77.36 49.15 69.50
QuIP 4.00 3.58 5.38 76.01 80.25 61.97 74.28 47.01 67.90
QuIP# 4.01 3.22 5.05 76.80 81.45 63.51 78.37 50.85 70.20
AQLM 3.82 3.21 5.03 76.32 80.90 63.69 77.61 50.34 69.77
E.1 Mixtral
We report the results for Mixtral(Jiang et al., 2024) MoE-type model for 3 and 4 bits in Table 11. In the 4 bit case, performance of QuIP# and AQLM are very similar across all metrics and close to uncompressed FP16 model.
Table 11: Evaluation of quantized Mixtral (Jiang et al., 2024) models for 3 and 4 bits per parameter. The table reports perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
3-bit 16.00 3.46 5.02 75.45 82.37 64.65 83.38 55.80 72.33
AQLM 3.02 3.8 5.18 74.74 81.77 63.22 82.28 52.30 70.86
4-bit 16.00 3.46 5.02 75.45 82.37 64.65 83.38 55.80 72.33
AQLM 3.915 3.57 5.07 74.82 81.99 64.23 83.12 54.61 71.75
QuIP# 4.000 3.60 5.08 76.56 81.99 63.92 82.62 54.78 71.97
E.2 Llama-2
We show results for 4 bit quantization of the Llama 2 models in Table 10. We can see that AQLM outperforms other methods in terms of perplexity and has the best or close to the best results. We also report results of perplexity for our quantized 2x8 codebooks models in Table 12.
Table 12: Wikitext2 and C4 PPL for quantized Llama 2 models for 2x8 codebooks models.
Name Wiki2normal-$\downarrow$ C4normal-$\downarrow$
7B 7.98 10.38
70B 4.83 6.66
Table 13: Evaluation of quantized Mistral7B (Jiang et al., 2023) models for 2, 3 and 4 bits per parameter: perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
2-bit 16.00 4.77 5.71 73.64 80.47 61.15 78.87 49.23 68.67
AQLM 2.01 6.32 6.93 68.75 76.01 52.13 73.65 40.44 62.17
QuIP# 2.01 6.02 6.84 69.30 76.71 52.95 72.14 39.76 62.20
3-bit 16.00 4.77 5.71 73.64 80.47 61.15 78.87 49.23 68.67
AQLM 3.04 5.07 5.97 72.69 80.14 59.31 77.61 46.67 67.28
4-bit 16.00 4.77 5.71 73.64 80.47 61.15 78.87 49.23 68.67
AQLM 4.02 4.89 5.81 73.80 79.71 60.27 77.86 48.21 67.97
QuIP# 4.01 4.85 5.79 73.95 80.41 60.62 78.96 49.40 68.67
E.3 Mistral
Finally we evaluate AQLM and QuIP# quantization on Mistral7b (Jiang et al., 2023) model for 3 and 4 bits in Table 13. In 2bits, QuIP# slightly outperform AQLM on most benchmarks. And for 4bits setup results are very close across the board.
Appendix F Pareto optimality
We visualize WikiText perplexity of Llama-2 7B,13B,70B models quantized with AQLM and QuIP# as plotted against quantized weight size in bytes and report it in Figure 5. Our method outperforms QuIP# in terms of perplexity in WikiText2 across all model sizes.
Additionally, in Figure 6, we show perplexity on WikiText for AQLM method against size of quantized parameters. We can notice that starting around 3.7 GiB of quantized weights, which correspond to 2.5 bits compression on Llama-2 13B model, it is more advantageous to compress 13B model rather 7B model at the same model size in bytes.
[Uncaptioned image] Figure 5: Comparison of AQLM relative to QuIP# on LLAMA-2 7B, 13B, and 70B models.
[Uncaptioned image] Figure 6: Model optimality for AQLM on LLAMA 2 7, 13, and 70B models.
Appendix G Estimating model size
In this section, we describe how to estimate the size of the quantized model for a given codebook configuration. The total cost of storing quantized weight comprises the codebooks, codes and per-unit scales. Specifically for a weight with input dimension $d_{in}$, output dimension $d_{out}$, group size $g$, $M$ codebooks corresponding to $B$-bit codes, the total amount of memory required is (assuming that codebooks and scales are stored in half precision):
  • codebooks: $g\cdot M\cdot 2^{B}\cdot 16$
  • codes: $d_{out}\cdot(d_{in}/g)\cdot B$
  • scales: $d_{out}\cdot 16$
Therefore, the average bits per parameter can be computed as follows:
[…]
For example, for mlp.gate_proj layer of Llama 2 70B model with $d_{in}=8192$, $d_{out}=28672$, quantization with group size $8$, two $8$-bit codebooks the formula above yields $2.002$ bits per parameter. Typically, storage cost is dominated by the codes, whereas codebooks and scales induce small memory overhead.
Appendix H End-to-End Inference Speed
Table 14: Text generation speed benchmark.
Llama 2 7B 13B 70B
Inference on Nvidia RTX 3090 GPU, tok/s
Original (float16) 41.51 26.76 5.66
AQLM (Table 1) 32.22 25.04 5.65
AQLM (2×2$\times$2 ×8-bit) 32.64 25.35 6.03
Inference on Intel i9 CPU, 8 cores, tok/s
Original (float32) 3.106 1.596 0.297
AQLM (2×2$\times$2 ×8-bit) 6.961 4.180 0.966
AQLM (4×4$\times$4 ×8-bit) 6.837 4.004 0.948
AQLM (8×8$\times$8 ×8-bit) 5.319 3.193 0.775
For quantized Llama-2 models, the same setup as in Section 4.3, we measure the time it takes to generate 128 tokens from scratch with batch size 1, and report the average number of generated tokens per second on a single 24GB RTX 3090 GPU, as well as Intel i9 CPU, in Table 14.
Appendix I Codebook and codes distribution
The proposed AQLM quantization method allows for large freedom in the choice of quantization lattice and ability to represent different weight distribution. To understand how do the learned codes and codebooks look like, we visualize the distribution of codes (how frequently given codebook vector is chosen) and the learned codebooks. Below on Figure 7 we provide a histogram of leaned codes and two leading principal components for a specific layer. One can observe that codes are assigned uniformly across codebook vectors and codebook vectors are concentrated in some ball. This pattern is pertinent to all linear projections inside transformer blocks.
Figure 7: Visualization of learned codes and codebooks in layers.31.mlp.down_proj linear projection.
(Left) Codes distribution. (Right) Two leading principal components of codebook.